# Signatures of spinon dynamics and phase structure of dipolar-octupolar quantum spin ices in two-dimensional coherent spectroscopy

### v1

@@ Abstract
The two-band dispersion of spinons in the $\pi$-flux phase has already been identified as a key signature distinguishing it from the 00-flux regime in linear response Lee et al. (2012); Desrochers and Kim (2024b), and we find that spinon transitions between these two bands produce features in the 2DCS response that clearly distinguish the two phases. Figure 3 presents the 2DCS responses for both phases calculated analytically using GMFT. For the 0-flux phase, a single rephasing streak is obtained, whilst the $\pi$-flux response resembles a cluster of isolated peaks. The reason for this stark contrast is two-fold. First, each of the two distinct spinon bands in the $\pi$-flux phase are relatively flat, leading to a density of states with two well defined maxima, and hence relatively narrow bandwidth optical responses. Second, the probe operator has non-zero matrix elements to excite a spinon from one band to another, and these inter-band transitions result in the additional peaks off of the main rephasing diagonal. Details of this calculation are given in the supplementary materialsup .
@@
Discussion- The 2DCS responses of spinons in dipolar-octupolar QSI are found to be sensitive to both temperature regime and phase of the system. Sharp responses reflecting the presence of fractionalized spinon excitations are predicted for the lowest $T$, and these responses are demonstrated to clearly distinguish between the spinon dynamics found in the 0-flux and $\pi$-flux phases of dipolar-octupolar spin ice. Conversely, in the intermediate temperature regime, a broad non-linear response is observed indicative of heavily constrained spinon motion in the presence of an incoherent spin background.
@@
Whilst this letter has focused on the example of dipolar-octupolar pyrochlores, we hypothesise that the appearance of broad signatures in 2DCS where fractionalised excitations move across an incoherent spin background may occur more generally in similar frustrated models, presenting an obstacle to observing sharp signatures of fractionalisation at all but the lowest temperatures. On the other hand, by measuring the 2DCS signal as a function of temperature one could use the crossover between these regimes to demonstrate the onset of a quantum coherent low-temperature state. The GMFT calculation however does not include the effects of quantum fluctuations in the gauge field, which will be necessary to understand the cross-over between these two regimes.
@@
A natural question is then whether the clear distinctions between the different spinon dynamical regimes observable in 2DCS can be observed more directly in the linear optical response to a uniform magnetic field, which is more simply accessed experimentally. In Fig. 4 we present the linear responses for all phases and regimes considered above. The linear response varies between each regime, and in particular the existence of multiple maxima in the $\pi$-flux response distinguishes it from the $T=0$ 00-flux response, as already noted in Ref. Desrochers and Kim (2024a). However, the stark difference between the behaviour of the response in the intermediate temperature regime and the low temperature regime is entirely absent; the intermediate temperature response and the low temperature 00-flux response are not qualitatively different, despite very different underlying spinon dynamics. By contrast, the existence of broad features in the intermediate temperature 2DCS response clearly distinguishes this regime from the low-temperature regime.
@@
To summarise, we find that two-dimensional coherent spectroscopy is a suitable experimental setting to probe the dynamics of fractionalised excitations in candidate quantum spin ice materials, providing clear signatures distinguishing the regime where fractionalised quasiparticles move ballistically on the pyrochlore lattice at the lowest temperatures from that where their motion is heavily constrained by an incoherent spin background. Further, the appearance of response streaks away from the main rephasing diagonal in 2DCS offers a direct signature of the transition of quasiparticles between bands, which for dipolar-octupolar quantum spin ices allows one to distinguish experimentally between a 00-flux or $\pi$-flux QSI ground states. We believe that this further demonstrates the potential utility of 2DCS as a probe of the dynamics of fractionalised excitations in quantum spin ice candidate materials, and other systems potentially hosting exotic quantum phases of matter.
Acknowledgements

### v2

@@ Abstract
The two-band dispersion of spinons in the $\pi$-flux phase has already been identified as a signature distinguishing it from the 00-flux regime in linear response [29, 37], and we find that spinon transitions between these two bands produce features in the 2DCS response that clearly distinguish the two phases. Figure 3 presents the 2DCS responses for both phases calculated within GMFT using the physical field coupling operator $M(t)=\mathbf{H}\cdot\mathbf{z}_{i}/|\mathbf{H}|\ S^{z}_{i}(t)$. For the 0-flux phase, a single rephasing streak is obtained, whilst the $\pi$-flux response resembles a cluster of isolated peaks. The reason for this stark contrast is two-fold. First, each of the two distinct spinon bands in the $\pi$-flux phase are relatively flat leading to a density of states with two well defined maxima, and hence relatively narrow bandwidth optical responses. Second, the probe operator has non-zero matrix elements to excite a spinon from one band to another, and these inter-band transitions result in the additional peaks off of the main rephasing diagonal.
@@
Discussion – The 2DCS responses of spinons in dipolar-octupolar QSI are found to be sensitive to both temperature regime and phase of the system. Sharp responses reflecting the presence of fractionalized spinon excitations are predicted for the lowest $T$, and these responses are demonstrated to clearly distinguish between the spinon dynamics found in the 0-flux and $\pi$-flux phases of dipolar-octupolar spin ice. Conversely, in the intermediate temperature regime, a broad non-linear response is observed indicative of heavily constrained spinon motion in the presence of an incoherent spin background.
@@
Whilst this letter has focused on the example of dipolar-octupolar pyrochlores, we hypothesize that the appearance of broad signatures in 2DCS where fractionalised excitations move across an incoherent spin background may occur more generally in similar frustrated models, presenting an obstacle to observing sharp signatures of fractionalisation at all but the lowest temperatures. On the other hand, by measuring the 2DCS signal as a function of temperature one could use the crossover between these regimes to demonstrate the onset of a quantum coherent low-temperature state. The GMFT calculation however does not include the effects of quantum fluctuations in the gauge field or interactions between the spinons and photon modes, which will be necessary to understand the cross-over between these two regimes.
@@
A natural question is then whether the clear distinctions between the spinon dynamical regimes observable in 2DCS can be seen directly in the linear optical response to a uniform magnetic field, which is more simply accessed experimentally. In Fig. 4 we present the linear responses for all phases and regimes considered above. The linear response varies between each regime, and in particular the existence of multiple maxima in the $\pi$-flux response distinguishes it from the $T=0$ 00-flux response, as already noted in Ref. [31]. However, the stark difference between the behavior of the response in the intermediate temperature regime and the low temperature regime is entirely absent; the intermediate temperature response and the low temperature 00-flux response are not qualitatively different, despite very different underlying spinon dynamics. By contrast, the existence of broad features in the intermediate temperature 2DCS response clearly distinguishes this regime from the low-temperature regime.
@@
To summarize, we find that 2DCS is a suitable experimental setting to probe the dynamics of fractionalised excitations in candidate quantum spin ice materials, providing clear signatures distinguishing the regime where fractionalised quasiparticles move ballistically on the pyrochlore lattice at the lowest temperatures from that where their motion is heavily constrained by an incoherent spin background. Further, the appearance of response streaks away from the main rephasing diagonal in 2DCS offers a direct signature of the transition of quasiparticles between bands, which for dipolar-octupolar quantum spin ices allows one to distinguish experimentally between a 00-flux or $\pi$-flux QSI ground states. We believe that this further demonstrates the potential utility of 2DCS as a probe of the dynamics of fractionalised excitations in quantum spin ice candidate materials, and other systems potentially hosting exotic quantum phases of matter.
Acknowledgements.